### v2

# Adversarial Training: Enhancing Out-of-Distribution Generalization for Learning Wireless Resource Allocation

Abstract
[…]
I Introduction
Deep learning is widely used for optimizing wireless resource allocation, by designing deep neural networks (DNNs) to learn the mapping from environment parameters (say channels) into optimized variables [1]. For wireless resource optimization with closed-form expressions of objective function and constraints, DNNs can be trained in an unsupervised manner. As a consequence, the labels obtained by solving computation-intensive numerical algorithms are no longer required, and constraint satisfaction can be controlled [1, 2].
The majority of existing works have assumed that training and test data are identically distributed. However, when the DNNs are tested on data drawn from unseen distributions, their out-of-distribution (OOD) performance often degrades obviously. In wireless communications, the distribution of environment parameters, especially channels, may vary rapidly due to many factors such as the changed locations of users and scatterers [3]. As a result, the distributions of training and test data may differ. A strong OOD generalizability enables DNNs to perform well across scenarios without further adaptation. This reduces the computational demands at wireless edge and aligns with the vision of green artificial intelligence (AI).
A few research efforts have attempted to enhance the OOD generalization of deep learning for wireless applications.
A data generation method was proposed in [4] to synthesize channel samples, based on an observation that DNNs trained on channels with proper power delay profiles are generalized well to other distributions when learning channel estimation. A data augmentation method was proposed in [5] after finding the major differences among channel distributions that are critical for channel feedback. These methods leverage the knowledge of certain tasks and are hard to be applied to other tasks.
It was reported that deep unfolding networks can achieve acceptable performance on unseen distributions when optimizing hybrid precoding[6] and power control[7], even without dedicated design for OOD generalizability. It was shown in [8] that the hyper-networks generating weights for the unfolding networks can achieve preferable OOD performance for signal detection when trained on multiple distributions. However, deep unfolding networks are problem- and algorithm-specific, which restricts their applicability.
Feature extractor (which is a DNN) was designed to capture the domain-invariant feature that does not change across distributions from each channel input, for learning reconfigurable intelligent surface configuration selection in [9] and indoor localization in [10]. The extracted feature was then fed into another DNN to produce the desired output. Nonetheless, the domain-invariant feature may not exist for all wireless problems, say precoding [11]. Meanwhile, samples from multiple distributions are required for training the feature extractor, which may not always be available in practice.
DNNs can be fine-tuned with transfer learning or meta-learning. Yet such adaptations are often infeasible in dynamic wireless environments [3] where the channel statistics change in tens of seconds [12]. This is because collecting and training with up-to-date samples are time-consuming and limited by the computing resource at wireless edge. In [13], a DNN for learning hybrid precoding was trained by meta-learning and evaluated for OOD generalization without adaptation, whose performance is only about half of that with sufficiently fine-tuning.
To achieve good OOD performance, distributionally robust optimization (DRO) has been introduced into machine learning [14]. In order to be well-generalized across a range of distributions, DRO optimizes the trainable weights of a DNN to perform better on the worst-case distribution where the DNN deteriorates most severely. Unfortunately, DRO problem is hard to be solved since it involves the optimization of both the weights and the worst-case distribution [14].
Adversarial training (AT) methods, originally proposed to defend against adversarial attacks, have been used as an easy-to-implement alternative to DRO [15, 16]. Specifically, it was proved that the objective of DRO is upper bounded by the objective of AT asymptotically (say Corollary 2 in [15], Theorems 1 and 2 in [16]). Consequently, DNNs trained by AT can also exhibit favorable OOD generalizability [16].
Both DRO and AT have been investigated only in the context of supervised learning, and can not be applied to unsupervised learning. The detection and defense for adversarial attacks has also been studied in supervised learning-based wireless AI [17]. However, the viability of using AT to enhance OOD generalization has not been explored for wireless communications.
In this paper, we resort to AT for improving the OOD generalizability of the DNNs that are trained unsupervisedly for optimizing resource allocation. The major contributions are summarized as follows. (1) To find the worst-case distribution in unsupervised learning, we reformulate DRO and AT problems, enabling them to be aware of the deterioration of DNNs. (2) We propose a one-step gradient ascent algorithm, referred to as unsupervised AT (UAT), to solve the reformulated AT problem. (3) We evaluate the OOD generalization gain from the UAT on diverse channel distributions by training several DNNs only on a single channel distribution, for optimizing hybrid precoding that maximizes sum-rate (SR) under quality of service (QoS) constraints.
The proposed UAT has three advantages: (1) It does not require specific expert knowledge to generate or augment data. (2) It does not need to change the architectures of DNNs, in contrast to deep unfolding and domain-invariant feature extraction. (3) It does not rely on data diversity, eliminating the need to access or distinguish samples from different distributions, in contrast to meta-learning, training hyper-networks or feature extractors.
II Preliminary
We first recap the DRO and AT in supervised learning.
Denote $\bf x$} as the input and $\bf y$$^{*}$ as the label for supervised learning, where the pair $({\bf x},{\bf y}^{*})$ follows a given distribution (denoted by $\mathcal{D}$) during training. A DNN denoted by $f_{\bf y}({\bf x};{\bm{\theta}})$ can produce an output $\bf y$} for each $\bf x$}, where the trainable weight $\bm{\theta}$ is optimized through a loss function $L({\bf y},{\bf y}^{*})$ to encourage the output $\bf y$} to approach the label $\bf y$$^{*}$.
To obtain satisfactory OOD performance on unknown distributions, DRO minimizes the expected loss on the worst-case distribution (denoted by $\tilde{\mathcal{D}}$) that incurs the maximal loss, which is formulated as [14]
$\min_{\bm{\theta}}\max_{\tilde{\mathcal{D}}\in\mathbb{D}}\mathbb{E}_{({\bf x},{\bf y}^{*})\sim\tilde{\mathcal{D}}}\left[L(f_{\bf y}({\bf x};{\bm{\theta}}),{\bf y}^{*})\right],$ (1)
where $\mathbb{D}=\{\mathcal{D}^{\prime}|W(\mathcal{D}^{\prime},\mathcal{D})\leq\varepsilon\}$ is the set of distributions around $\mathcal{D}$, $W(\cdot,\cdot)$ denotes the Wasserstein distance between two distributions, and $\varepsilon$ is a threshold.
Although solving the problem in (1) enables the DNN to achieve minimal loss over all distributions in $\mathbb{D}$, which improves its OOD generalizability, the impossibility of sampling from the unsolved $\tilde{\mathcal{D}}$ renders the problem intractable [14].
AT methods find adversarial examples and adjust the trainable weights of DNNs with these examples to prevent possible attacks. Specifically, AT problem minimizes the expected loss over the most maliciously perturbed inputs (denoted by $\tilde{\bf x}$) [18], which is formulated as follows
$\min_{\bm{\theta}}\mathbb{E}_{({\bf x},{\bf y}^{*})\sim\mathcal{D}}\left[\max_{\tilde{\bf x}\in B_{\epsilon}({\bf x})}L(f_{\bf y}(\tilde{\bf x};{\bm{\theta}}),{\bf y}^{*})\right],$ (2)
where $B_{\epsilon}({\bf x})=\{{\bf x}+{\bm{\delta}}\,|\left\|{\bm{\delta}}\right\|_{2}\leq\epsilon\}$ is the set of allowed perturbations around $\bf x$}, and $\epsilon$ is the perturbation range.
Differing from (1), the “optimization variable” $\tilde{\bf x}$ in (2) is a sample rather than a distribution, which is referred to as the adversarial example of $\bf x$}. Since the expectation in (2) can be approximated by averaging over samples drawn from $\mathcal{D}$, this problem can be solved by alternating the following two steps [18]: (i) Find $\tilde{\bf x}$: It can be obtained using gradient ascent methods, guided by the partial derivative of loss function with respect to $\bf x$}. (ii) Update $\bm{\theta}$}: It can be optimized using back-propagation, with the loss averaged over all adversarial examples.
III Adversarial Training in Unsupervised Learning
In this section, we formulate DRO and AT for unsupervised learning, and propose an algorithm for solving the AT problem.
III-A Problem Reformulation
With environment parameter $\bf x$}, resource allocation $\bf y$ can be optimized from the following problem
$\min_{\bf y}~F({\bf x},{\bf y})~~~{\rm s.t.}~G_{i}({\bf x},{\bf y})\leq 0,i=1,\cdots,N_{C}.$ (3)
The optimal solutions for every value of $\bf x$} constitute a function ${\bf y}^{*}=f_{{\bf y}^{*}}({\bf x})$. When a DNN ${\bf y}=f_{\bf y}({\bf x};{\bm{\theta}})$ is trained to learn this function, $\bf x$} follows the distribution $\mathcal{D}$.
By introducing a multiplier network (denoted by $f_{\bm{\lambda}}({\bf x};{\bm{\xi}})$), primal-dual learning (PDL) [2, 1] can be adopted to train $f_{\bf y}(\cdot\,;{\bm{\theta}})$ and $f_{\bm{\lambda}}(\cdot\,;{\bm{\xi}})$ from the following problem
$\max_{\bm{\xi}}\min_{\bm{\theta}}\mathbb{E}_{{\bf x}\sim\mathcal{D}}[\mathcal{L}({\bf x};{\bm{\theta}},{\bm{\xi}})],$ (4)
where $\mathcal{L}({\bf x};{\bm{\theta}},{\bm{\xi}})$ is the Lagrangian function defined as
$\mathcal{L}({\bf x};{\bm{\theta}},{\bm{\xi}})\!=\!F({\bf x},f_{\bf y}({\bf x};{\bm{\theta}}))+\!\sum_{i=1}^{N_{C}}\!f_{\lambda_{i}}\!({\bf x};{\bm{\xi}})G_{i}({\bf x},f_{\bf y}({\bf x};{\bm{\theta}})),\!$ (5)
and $f_{\lambda_{i}}({\bf x};{\bm{\xi}})$ represents the $i$th element of the output $\lambda_{1},\cdots,\lambda_{N_{C}}]^{T}=f_{\bm{\lambda}}({\bf x};{\bm{\xi}})$.
After being well-trained, the resource allocation variable provided by $f_{\bf y}(\cdot\,;{\bm{\theta}})$ can achieve near-optimal objective value in (3) and satisfy the constraints with high probability [2, 1].
To improve OOD performance, we should introduce a cost function [^1] In supervised learning, the cost function is the same as the loss function, since the loss can indicate the performance of DNNs. to find the worst-case distribution for DRO, which measures how worse the DNN performs on a distribution. In this paper, we focus on pursuing better objective on unseen distributions, and thus select $F({\bf x},{\bf y})$ as the cost function. [^2] If one seeks to further guarantee the constraints on unseen distributions, $F({\bf x},{\bf y})+\sum_{i=1}^{N_{C}}\max(0,G_{i}({\bf x},{\bf y}))$ can be used as the cost function, and the subsequent analysis is similar.
Analogous to (1), one can optimize $\bm{\theta}$} on the distribution $\tilde{\mathcal{D}}$ with maximal cost, which is the solution of $\max_{\tilde{\mathcal{D}}\in\mathbb{D}}\mathbb{E}_{{\bf x}\sim\tilde{\mathcal{D}}}\left[F({\bf x},f_{\bf y}({\bf x};{\bm{\theta}}))\right]$. However, $\tilde{\mathcal{D}}$ obtained in this way is NOT the worst-case distribution, because this objective cannot differentiate the cost increase caused by the higher optimal objective value and by the degraded generalization performance of $f_{\bf y}(\cdot\,;{\bm{\theta}})$ on the distributions in $\mathbb{D}$. The former is unrelated to generalizability. For example, for SR-maximization resource allocation, a decrease in channel strength inevitably increases the cost (i.e., reduces SR), but a DNN may not perform worse on low-strength channels. Optimizing the DNN on such channels is not beneficial for OOD generalization.
To address this issue, [^3] This issue does not arise in supervised learning, since the optimal value of the losses (such as cross-entropy, mean squared error) is always zero. we reformulate DRO for PDL as
$P_{\rm DRO}:~\max_{\bm{\xi}}\min_{\bm{\theta}}\mathbb{E}_{{\bf x}\sim\tilde{\mathcal{D}}}[\mathcal{L}({\bf x};{\bm{\theta}},{\bm{\xi}})]$ (6a)
${\rm s.t.}$~ $\tilde{\mathcal{D}}\!=\!\arg\max_{\mathcal{D}^{\prime}\in\mathbb{D}}\!\mathbb{E}_{{\bf x}\sim\mathcal{D}^{\prime}}\!\left[F({\bf x},f_{\bf y}({\bf x};{\bm{\theta}}))\!-\!F({\bf x},f_{{\bf y}^{*}}({\bf x}))\right]\!,\!\!\!$ (6b)
for training $f_{\bf y}(\cdot\,;{\bm{\theta}})$ and $f_{\bm{\lambda}}(\cdot\,;{\bm{\xi}})$ on the truely worst distribution, where the output of $f_{\bf y}(\cdot\,;{\bm{\theta}})$ yields the largest gap in cost from the optimal solution instead of the highest cost.
Inspired by the idea in [16], we adopt AT as a realizable alternative to the DRO for enhancing OOD generalizability performance. By replacing $\tilde{\mathcal{D}}$ by $\tilde{\bf x}$, the AT problem for PDL is reformulated as
$P_{\rm AT}:~\max_{\bm{\xi}}\min_{\bm{\theta}}\mathbb{E}_{{\bf x}\sim\mathcal{D}}[\mathcal{L}(\tilde{\bf x};{\bm{\theta}},{\bm{\xi}})]$ (7a)
${\rm s.t.}$~ $\tilde{\bf x}=\arg\!\!\!\!\max_{{\bf x}^{\prime}\in B_{\epsilon}({\bf x})}\!\!\left[F({\bf x}^{\prime},f_{\bf y}({\bf x}^{\prime};{\bm{\theta}}))-F({\bf x}^{\prime},f_{{\bf y}^{*}}({\bf x}^{\prime}))\right].\!$ (7b)
III-B Unsupervised Adversarial Training (UAT) Algorithm
To solve $P_{\rm AT}$, i.e., to train $f_{\bf y}(\cdot\,;{\bm{\theta}})$ and $f_{\bm{\lambda}}(\cdot\,;{\bm{\xi}})$ with found adversarial examples, (7b) and (7a) can be alternately optimized. To find accurate adversarial examples, both DNNs need to first be trained with in-distribution (ID) samples, namely, original samples provided in the training set.
In what follows, we show how to find the adversarial example $\tilde{\bf x}={\bf x}+{\bm{\delta}}$ for every original sample $\bf x$} from (7b).
From the definition of $B_{\epsilon}({\bf x})$ in (2), (7b) can be transformed into the following problem to optimize $\bm{\delta}$}
$\max_{\left\|{\bm{\delta}}\right\|_{2}\leq\epsilon}\!F({\bf x}+{\bm{\delta}},f_{\bf y}({\bf x}+{\bm{\delta}};{\bm{\theta}}))-F({\bf x}+{\bm{\delta}},f_{{\bf y}^{*}}({\bf x}+{\bm{\delta}})).\!\!$ (8)
To find $\bm{\delta}$} (and hence find $\tilde{\bf x}$) with low complexity, we employ a one-step gradient ascent method. By using the first-order Taylor polynomial, it can be seen that (8) has the same solution as the following problem when $\epsilon$ is very small,
$\max_{\left\|{\bm{\delta}}\right\|_{2}\leq\epsilon}\left(\frac{\partial F({\bf x},f_{\bf y}({\bf x};{\bm{\theta}}))}{\partial{\bf x}}-\frac{{\rm d}F({\bf x},f_{{\bf y}^{*}}({\bf x}))}{{\rm d}{\bf x}}\right)^{T}{\bm{\delta}}.$ (9)
The problem in (9) maximizes a linear objective over an Euclidean ball, whose optimal solution can be obtained as
${\bm{\delta}}=\epsilon\cdot\frac{\frac{\partial F({\bf x},f_{\bf y}({\bf x};{\bm{\theta}}))}{\partial{\bf x}}-\frac{{\rm d}F({\bf x},f_{{\bf y}^{*}}({\bf x}))}{{\rm d}{\bf x}}}{\left\|\frac{\partial F({\bf x},f_{\bf y}({\bf x};{\bm{\theta}}))}{\partial{\bf x}}-\frac{{\rm d}F({\bf x},f_{{\bf y}^{*}}({\bf x}))}{{\rm d}{\bf x}}\right\|_{2}}.$ (10)
To obtain the value of the numerator in (10), we rewrite its first term by using the chain rule as
$\frac{\partial F({\bf x},f_{\bf y}({\bf x};{\bm{\theta}}))}{\partial{\bf x}}=\frac{\partial F}{\partial{\bf x}}+\frac{\partial F}{\partial{\bf y}}\cdot\frac{\partial f_{\bf y}({\bf x};{\bm{\theta}})}{\partial{\bf x}},$ (11)
where the values of partial derivatives $\frac{\partial F}{\partial{\bf x}}$ and $\frac{\partial F}{\partial{\bf y}}$ can be computed using the value of $\bf y$} produced by $f_{\bf y}({\bf x};{\bm{\theta}})$, and $\frac{\partial f_{\bf y}({\bf x};{\bm{\theta}})}{\partial{\bf x}}$ can be obtained from the DNN due to its differentiability.
Obtaining the value of the second term of the numerator is challenging, due to the unknown function $f_{{\bf y}^{*}}(\cdot)$. To cope with this difficulty, we invoke the envelope theorem [19], which provides the gradient of the optimal objective from (3), with no need for the closed-form expression of $f_{{\bf y}^{*}}(\cdot)$, as
$\frac{{\rm d}F({\bf x},f_{{\bf y}^{*}}({\bf x}))}{{\rm d}{\bf x}}=\frac{\partial F}{\partial{\bf x}}+\lambda_{1}^{*}\frac{\partial G_{1}}{\partial{\bf x}}+\cdots+\lambda_{N_{C}}^{*}\frac{\partial G_{N_{C}}}{\partial{\bf x}},$ (12)
where $\lambda_{1}^{*},\cdots,\lambda_{N_{C}}^{*}$ are the optimal Lagrange multipliers for the $N_{C}$ constraints in (3), which depend on $\bf x$}.
Unlike (11), the values of partial derivatives in (12) are computed with the value of optimal solution $\bf y$$^{*}$. In practice, the value of $\bf y$$^{*}$ can be approximated by $\bf y$} (i.e., the output of $f_{\bf y}({\bf x};{\bm{\theta}})$), since $\bf x$} herein is an ID sample and $f_{\bf y}(\cdot;{\bm{\theta}})$ has been trained with all ID samples before UAT begins. Similarly, the values of $\lambda_{1}^{*},\cdots,\lambda_{N_{C}}^{*}$ can be approximated by the outputs of multiplier network $f_{\bm{\lambda}}({\bf x};{\bm{\xi}})$.
By combining (11) and (12), we have
$\frac{\partial F({\bf x},f_{\bf y}({\bf x};\!{\bm{\theta}})\!)\!}{\partial{\bf x}}-\frac{{\rm d}F({\bf x},f_{{\bf y}^{*}}\!({\bf x})\!)\!}{{\rm d}{\bf x}}\!=\!\frac{\partial F}{\partial{\bf y}}\cdot\frac{\partial f_{\bf y}}{\partial{\bf x}}-\!\sum_{i=1}^{N_{C}}\!f_{\lambda_{i}}\!\frac{\partial G_{i}}{\partial{\bf x}}.\!\!\!$ (13)
By substituting (13) into (10), the adversarial example of $\bf x$}, i.e., $\tilde{\bf x}={\bf x}+{\bm{\delta}}$, can be obtained as
$\tilde{\bf x}\!=\!{\bf x}\!+\epsilon\cdot\frac{\frac{\partial F}{\partial{\bf y}}\!\cdot\!\frac{\partial f_{\bf y}({\bf x};{\bm{\theta}})}{\partial{\bf x}}\!-\!\!\sum_{i=1}^{N_{C}}\!f_{\lambda_{i}}\!({\bf x};{\bm{\xi}})\frac{\partial G_{i}}{\partial{\bf x}}}{\left\|\!\frac{\partial F}{\partial{\bf y}}\!\cdot\!\frac{\partial f_{\bf y}({\bf x};{\bm{\theta}})}{\partial{\bf x}}\!-\!\!\sum_{i=1}^{N_{C}}\!f_{\lambda_{i}}\!({\bf x};{\bm{\xi}})\frac{\partial G_{i}}{\partial{\bf x}}\!\right\|_{2}\!\!\!}\!\triangleq\!{\bf x}\!+\epsilon\cdot d({\bf x}).\!\!$ (14)
Remark 1: The constraints in (3) for some wireless problems are imposed on $\bf y$} but not related to $\bf x$}. For instance, power constraint is independent of channels. Such constraints can be ensured via projection methods, and render the multiplier network $f_{\bm{\lambda}}({\bf x};{\bm{\xi}})$ unnecessary. For these problems, $\frac{\partial G_{i}}{\partial{\bf x}}=0$ in (12). By contrast, for the problems with complex constraints (say the QoS constraints) that are related to $\bf x$} (say channels), $f_{\bm{\lambda}}({\bf x};{\bm{\xi}})$ is required and the values of corresponding partial derivatives $\frac{\partial G_{i}}{\partial{\bf x}}$ in (12) ought to be computed.
With a batch of adversarial examples, (7a) can be optimized by standard back-propagation algorithms (say Adam algorithm), where $\mathcal{L}(\cdot\,;{\bm{\theta}},{\bm{\xi}})$ is computed using the adversarial examples rather than the original samples.
During UAT, (7b) and (7a) are alternately solved for $T$ iterations. In the first iteration, the adversarial examples are generated from the original samples using (14). In the subsequent iterations, the adversarial examples are generated from those obtained in the previous iteration. In this way, the range of perturbations can be expanded for facilitating OOD generalization, while in each iteration $\epsilon$ can be set as a small value for the Taylor approximation being accurate.
[…]
Algorithm 1 Unsupervised Adversarial Training (UAT)
1:Input: number of epochs $E$, number of batches $B$, batch-size $N_{B}$, samples $\{{\bf x}_{n}\}_{n=1}^{BN_{B}}$, initial value of weight $\bm{\theta}$}, learning rate $\eta_{1},\eta_{2}$, the epoch at which UAT begins $E_{A}$, number of adversarial iterations $T$, perturbation range $\epsilon$
2:* initial weight of multiplier network $\bm{\xi}$}
3:for $epoch=1,\dots,E$ do
4:  for $batch=1,\dots,B$ do
5:   $\mathcal{N}=\{1+N_{B}(batch-1),\cdots,N_{B}(batch)\}$
6:   $\mathcal{L}=\frac{1}{N_{B}}\sum_{n\in\mathcal{N}}F({\bf x}_{n},f_{\bf y}({\bf x}_{n};{\bm{\theta}}))$
7:   $\mathcal{L}\!\leftarrow\!\mathcal{L}\!+\frac{1}{N_{B}}\!\!\sum_{n}\!\!\sum_{i=1}^{N_{C}}\!f_{\lambda_{i}}({\bf x}_{n};{\bm{\xi}})G_{i}({\bf x}_{n},f_{\bf y}({\bf x}_{n};{\bm{\theta}}))$
8:   ${\bm{\theta}}\leftarrow{\bm{\theta}}-\eta_{1}\frac{\partial\mathcal{L}}{\partial{\bm{\theta}}}$
9:   ${\bm{\xi}}\leftarrow{\bm{\xi}}+\eta_{2}\frac{\partial\mathcal{L}}{\partial{\bm{\xi}}}$
10:   if $epoch\geq E_{A}$ then $\triangleright$ UAT begins
11:     $\tilde{\bf x}_{n}^{0}={\bf x}_{n},~~\forall n\in\mathcal{N}$
12:     for $t=1,\cdots,T$ do
13:       $\tilde{\bf x}_{n}^{t}=\tilde{\bf x}_{n}^{t-1}\!+\epsilon\cdot d(\tilde{\bf x}_{n}^{t-1}),\!\forall n\!\in\!\mathcal{N}$ $\triangleright$ see (14)
14:       $\mathcal{L}=\frac{1}{N_{B}}\sum_{n\in\mathcal{N}}F(\tilde{\bf x}_{n}^{t},f_{\bf y}(\tilde{\bf x}_{n}^{t};{\bm{\theta}}))$
15:       $\mathcal{L}\!\!\leftarrow\!\!\mathcal{L}\!+\!\frac{1}{N_{B}}\!\!\sum_{n}\!\!\sum_{i=\!1}^{\!N_{C}}\!\!f_{\!\lambda_{i}}\!(\!\tilde{\bf x}_{n}^{t};\!{\bm{\xi}}\!)G_{i}\!(\tilde{\bf x}_{n}^{t},\!f_{\bf y}(\tilde{\bf x}_{n}^{t};\!{\bm{\theta}})\!)$
16:       ${\bm{\theta}}\leftarrow{\bm{\theta}}-\eta_{1}\frac{\partial\mathcal{L}}{\partial{\bm{\theta}}}$ $\triangleright$ via back-propagation
17:       ${\bm{\xi}}\leftarrow{\bm{\xi}}+\eta_{2}\frac{\partial\mathcal{L}}{\partial{\bm{\xi}}}$ $\triangleright$ via back-propagation
18:Output: trained weight $\bm{\theta}$}
IV Case Study: Optimizing Hybrid Precoding
In this section, we apply UAT for optimizing QoS-constrained hybrid precoding.
Consider a base station (BS) equipped with $N_{T}$ antennas and $N_{RF}$ radio frequency (RF) chains, which serves $K$ single-antenna users. The analog and baseband precoders can be jointly optimized to maximize SR as follows [20],
$\max_{{\bf W}_{RF},{\bf W}_{BB}}~~\textstyle\sum_{k=1}^{K}R_{k}$ (15a)
${\rm s.t.}~~\left\|{\bf W}_{RF}{\bf W}_{BB}\right\|^{2}_{F}=P_{tot},$ (15b)
$~~|\left({\bf W}_{RF}\right)_{n,j}|=1,n=1,\cdots,N_{T},j=1,\cdots,N_{RF},$ (15c)
$~~R_{k}\geq\gamma_{k},k=1,\cdots,K,$ (15d)
where $R_{k}\!=\!\log_{2}\!\left(\!1\!+\!\frac{|{\bf h}^{H}_{k}{\bf W}_{RF}{\bf w}_{BB_{k}}|^{2}}{\sum_{i=1,i\neq k}^{K}\!|{\bf h}^{H}_{k}{\bf W}_{RF}{\bf w}_{BB_{i}}|^{2}\!+\sigma^{2}}\!\!\right)$, ${\bf h}_{k}\in\mathbb{C}^{N_{T}}$, and $\gamma_{k}$ are the achievable rate, channel vector, and data-rate requirement of the $k$th user, respectively, ${\bf W}_{RF}\in\mathbb{C}^{N_{T}\times N_{RF}}$ is the analog precoder, ${\bf W}_{BB}=[{\bf w}_{BB_{1}},\cdots,{\bf w}_{BB_{K}}]\in\mathbb{C}^{N_{RF}\times K}$ is the baseband precoder, $P_{tot}$ is the total power, and $\sigma^{2}$ is the noise power. (15b) is the power constraint, (15c) is the constant modulus constraint for the analog precoder, and (15d) is the QoS constraint for every user.
[…]
When applying UAT, we only need to generate adversarial examples for ${\bf x}={\bf H}/\!\left\|{\bf H}\right\|_{F}$, since SNRS\!N\!R is a scalar with finite range that does not need to be generalized. To ensure that the adversarial examples are normalized, we modify (14) as
$\tilde{\bf x}=\frac{{\bf x}+\epsilon\cdot d({\bf x})}{\left\|{\bf x}+\epsilon\cdot d({\bf x})\right\|_{2}}.$ (16)
V Simulation Results
In this section, we evaluate the OOD performance of DNNs trained by the proposed UAT for the hybrid precoding task.
We use the following two metrics to measure performance. Average SR (ASR) is the averaged SR over a test set, where the SR of a sample is recorded as zero if the constraints are not satisfied. Violation Ratio (VR) is the proportion of samples in a test set that violate the constraints.
To obtain a benchmark for comparison, we simulate a difference-of-convex (DC) programming-based numerical algorithm [20] to optimize hybrid precoders. Yet different from [20], we set ${\bf W}_{RF}$ as the phases of the first $N_{RF}$ right singular vectors of $\bf H$}, instead of selecting from the array responses computed by the angles of departure (AoDs) of all propagation paths that is only applicable to sparse channels.
[…]
We compare UAT with two existing AT methods in [18, 16] that generate adversarial examples for supervised learning. Following the configurations in [18, 16], the number of descent steps is 7, step size is 2, and projection radius is 8. For UAT, the fine-tuned hyper-parameters are $E_{A}=50$, $T=5$, $\epsilon=0.5$. The PDL in [1] without any AT serves as a baseline.
Both $f_{\bf y}(\cdot\,;{\bm{\theta}})\!$ and $f_{\bm{\lambda}}(\cdot\,;{\bm{\xi}})\!$ are realized by three kinds of DNNs designed for hybrid precoding: convolutional neural networks (CNN) [22], three-dimensional graph neural network (3D-GNN) [21], and recursive GNN (RGNN) [23]. The fine-tuned hyper-parameters are as follows. The optimizer is Adam, $\eta_{1}=\eta_{2}=0.001$, $N_{B}=100$, batch normalization is used for CNN and 3D-GNN, activation function is ${\rm Tanh}(\cdot)$ for RGNN and ${\rm ReLU}(\cdot)$ for CNN and 3D-GNN. Other fine-tuned hyper-parameters are listed in Table I, where $3\times 3$ convolution kernels are used in the hidden layers of CNN followed by a fully-connected layer with 512 neurons.
TABLE I: Tuned Hyper-Parameters for DNNs
Kind of DNN $f_{\bf y}(\cdot~;{\bm{\theta}})$ $f_{\bm{\lambda}}(\cdot~;{\bm{\xi}})$
Number of hidden layers Number of channels in layers Number of hidden layers Number of channels in layers
CNN 5 [256, $\cdots$, 256, 16] 3 [128, 128, 16]
3D-GNN 6 [256, $\cdots$, 256] 4 [128, $\cdots$, 128]
RGNN 4 [64, 64, 64, 16] 3 [32, 32, 8]
Since the UAT does not depend on data diversity, these DNNs are trained with 100,000 samples generated from a single distribution, where we take Rayleigh fading channel as an example. In fact, they can also be trained on other channel distributions and yield similar results. Once trained, these DNNs are not fine-tuned, in order to examine their OOD generalization performance.
The performances of the DNNs trained with different methods are evaluated on several test sets, each consisting of 2,000 samples generated from unseen channel distributions (referred to as OOD test samples), as detailed in the sequel.
V-A Rician Fading Channel
The OOD test samples are generated from the Rician fading channel model: ${\bf h}_{k}=\sqrt{{\kappa}/({\kappa+1})}{\bf a}(\phi_{k})+\sqrt{{1}/({\kappa+1})}{\bf z}_{k}$, where ${\bf a}(\phi_{k})=[1,{\rm e}^{{\rm i}\pi\sin(\phi_{k})},\cdots,{\rm e}^{{\rm i}(N_{T}-1)\pi\sin(\phi_{k})}]^{T}$ is the array response of the line-of-sight (LOS) path with AoD $\phi_{k}\sim\mathcal{U}(0,2\pi)$, ${\bf z}_{k}\sim\mathcal{CN}({\bf 0},{\bf I}_{N_{T}})$, and $\kappa$ is the Rician factor.
In Fig. 1, we provide the performance of the RGNN trained by PDL or UAT (the results for the other two DNNs are similar and hence are not shown). It has legend PDL(ID) or UAT(ID) when tested in Rayleigh channel and it has legend PDL(OOD) or UAT(OOD) when tested in Rician channel. The Rician channel with $\kappa=10$ dB is a representative small-scale fading of LOS channels in UMa and UMi scenarios [24]. Fig. 1 (a) shows that RGNN trained by both methods can achieve comparable ASR to DC in Rayleigh channel. When tested in Rician channel, the RGNN trained by UAT attains a 4.7% higher ASR (normalized by DC) than the RGNN trained by PDL. Fig. 1 (b) shows that DC can satisfy QoS constraints for all samples on both channels. When tested in Rayleigh channel, RGNN trained by both methods exhibits extremely low VR (about 0.3%). When tested in Rician channels, the RGNN trained by UAT has a 4.9% lower VR than that trained by PDL, because the objective in (15a) and the constraints in (15d) are positively correlated, even though no constraint satisfaction is considered in the cost of our DRO formulation.
Besides, the inference time of RGNN is 15.48 ms, but DC requires 5498 ms on the same CPU, validating the advantage of learning-based approach for low decision latency.
(a) ASR performance
(b) VR performance
Figure 1: Performance of RGNN on ID test samples (Rayleigh) and OOD test samples (Rician, $\kappa=10$ dB), during training with PDL or UAT.
In Fig. 2, we compare the OOD generalization performance of RGNN and 3D-GNN trained by different methods. We can see that UAT improves OOD generalization by approximately 5% in both ASR and VR for the two GNNs when the distribution of test samples differs noticeably from Rayleigh channel (say $\kappa\geq 10$ dB). The ASR of UAT-trained DNNs still retains 85% of the ASR of DC even when $\kappa=20$ dB. The DNNs trained by the AT methods in [18, 16] exhibit marginal OOD generalization gains over PDL, due to not designed for unsupervised learning. Their performance degradation observed at $\kappa\leq 5$ dB (close to Rayleigh channels, i.e., ID samples) results from their overly broad perturbation range.
(a) ASR performance
(b) VR performance
Figure 2: OOD performance of two GNNs trained with different methods.
V-B Spatially Correlated Channel
The OOD test samples are generated from the following spatially correlated channel model [25]: ${\bf H}={\bf R}^{\frac{1}{2}}_{U}{\bf Z}{\bf R}^{\frac{1}{2}}_{A}$, where each element in $\bf Z$} is $z_{k,n}\sim\mathcal{CN}(0,1)$, ${\bf R}_{A}\in\mathbb{R}^{N_{T}\times N_{T}}$ is the antenna correlation matrix with element in the $i$th row and $j$th column being $\rho_{A}^{|i-j|}$, ${\bf R}_{U}\in\mathbb{R}^{K\times K}$ captures the inter-user correlation with diagonal elements being 1 and off-diagonal elements being $\rho_{U}$, and $\rho_{A},\rho_{U}\in[0,1]$.
TABLE II: ASR Performance (Relative to DC) in Correlated Channels
$\rho_{U}$ varies ($\rho_{A}=0$) $\rho_{A}$ varies ($\rho_{U}=0$)
0.2 0.4 0.6 0.8 0.2 0.4 0.6 0.8
DC (bps/Hz) 18.77 17.62 15.65 12.36 19.13 18.86 18.14 16.26
RGNN PDL 100.3% 99.2% 91.3% 72.6% 100.6% 99.1% 95.6% 78.3%
UAT 99.9% 98.8% 95.8% 87.4% 100.4% 100.3% 99.7% 98.6%
3D-GNN PDL 98.7% 97.3% 93.8% 79.2% 99.1% 98.5% 96.4% 84.9%
UAT 100.0% 99.4% 98.1% 93.0% 100.3% 99.9% 99.5% 95.5%
CNN PDL 82.2% 80.0% 74.2% 57.6% 82.6% 81.9% 80.2% 70.5%
UAT 92.5% 91.3% 89.2% 80.2% 92.8% 92.1% 91.3% 84.0%
[…]
V-C Sparse Channel and 3GPP Channel
Finally, the OOD test samples are generated from two millimeter-wave channel models, which are commonly used for hybrid precoding.
[…]
In Table III, we provide the OOD generalization performance of the DNNs trained by PDL or UAT when they are tested on the two kinds of channels. It is evident that UAT enhances the OOD generalizability of all three DNNs, under all considered channels for testing. The results in boldface indicate that 3D-GNN outperforms both RGNN and CNN in terms of OOD generalization.
TABLE III: ASR Performance (Relative to DC) in Sparse & 3GPP Channels
Sparse Channels Scenarios in TR 38.901
Number of paths, $L$ UMa NLOS UMa LOS UMi NLOS UMi LOS
5 4 3 2
DC (bps/Hz) 18.28 18.24 17.93 17.43 9.74 15.44 14.89 16.84
RGNN PDL 96.6% 95.0% 89.7% 80.2% 95.0% 88.1% 73.8% 82.6%
UAT 97.8% 96.1% 92.7% 84.6% 111.3% 100.5% 91.7% 96.1%
3D-GNN PDL 96.4% 95.7% 93.1% 86.0% 97.0% 91.5% 76.4% 87.2%
UAT 99.1% 98.4% 96.6% 90.2% 117.2% 103.9% 95.1% 100.3%
CNN PDL 77.4% 75.8% 70.5% 60.8% 95.9% 77.8% 69.4% 74.2%
UAT 87.9% 86.0% 80.8% 70.3% 102.6% 86.2% 77.3% 82.7%
VI Conclusions
In this paper, we reformulated AT problem for training OOD-generalizable DNNs in the unsupervised manner to optimize wireless resource allocation, and proposed an algorithm to solve the problem. We train three DNNs for learning QoS-aware hybrid precoding with the proposed UAT as a case study for a proof-of-concept demonstration. Simulation results showed that these DNNs trained by the proposed UAT can achieve superior OOD performance on Rician fading, spatially correlated, sparse, and 3GPP channels, despite being trained only on Rayleigh fading channels. In future works, we will evaluate the OOD generalization performance of the UAT over more realistic channels, such as real-world datasets.